OpenStudy (anonymous):

Can I please get help with this? Simplify the expression. Write the answer using scientific notation. (5 x 10^5)^-2 4.0 x 10^-11 0.04 x 10^-10 4.0 x 10^-12 -10 x 10^-10

OpenStudy (anonymous):

\[(5 \times 10^5)^{-2} \implies \frac{ 1 }{ (5 \times 10^5)^2 }\]

OpenStudy (anonymous):

this is the same as \[\frac{ 1 }{ 250,000,000,000 } \implies 4.0 \times 10^{-12}\] in scientific notation

OpenStudy (anonymous):

because \[(5 \times 10^5)^2 \implies 25 \times 10^{10}\] but yeah, I think you see the point haha.